OpenStudy (anonymous):

215=1/3pi*x^2*13.. solve for x

OpenStudy (anonymous):

are you allowed to use a calculator for this problem?

OpenStudy (anonymous):

yeah do you need a graph or the result of x?

OpenStudy (anonymous):

i just need to find the value of x... its for the volume of a cone

OpenStudy (anonymous):

\[X=\sqrt{645/13\pi}\]

OpenStudy (anonymous):

yay! i figured out the equations button!

OpenStudy (anonymous):

x is the radius of the cone

OpenStudy (anonymous):

thats a really gross equation because that's as simplified as it gets...

OpenStudy (anonymous):

this is correct

OpenStudy (anonymous):

so i dont do 13*pi/3=(13.61)x^2= 215/13.61=15.80

OpenStudy (anonymous):

\[so using x =\sqrt{645/13\pi }\] i got 12.48